OpenStudy (dehelloo):

Matrices question/ solving simultaneous equations. The ages of John cleese and Sudam Hussein add to 83.25 years. Cleese's age is 3.5 times that of Hussein's age. Write two simultaneous equations to show this. Find each of their ages, using matrices?

OpenStudy (anonymous):

What helps on these word problems is having variables for the names. :) So you can just shorthand it and start throwing away a lot of the words! I talk it out when I write it: "c plus h equal 83.25 years." --> c + h = 83.25 "c equals 3.5 times h" -->

OpenStudy (dehelloo):

Yeah. So would the two equations be c+h=83.25 c+ 3.5 x h= 83.25 I have a feeling that's wrong. I'm unsure how to construct two different equations.

OpenStudy (anonymous):

Oh! The 2nd equation is \[c = 3.5 h\]. There's no 83.25 there. You can then subtract c from both sides:\[c - c = 3.5h - c\] and then get as your second equation: \[0 = 3.5h - c\] so c+ h = 83.25 -c+3.5h = 0

OpenStudy (dehelloo):

Thanks mate